Reports 1-1 of 1 Clear search Modify search
AdV-COM (AdV commissioning (1st part) )
Allocca, Chiummo, Ruggi, Magazzu' - 2:09 Wednesday 23 January 2019 (44486) Print this report
ITF recovery

Tonight's shift was dedicated to the ITF recovery. 

We closed the PR angular loop in full bandwith in LN1, while the BS angular loop is only closed in drift control (put "1" instead of "2" in the BS.ramp_attribute("AATX(Y)_FLAG") in the ITF_LOCK.py) and we also slightly retuned the offsets.

The OMC1 lock didn't give much problems, but it was not possible to achieve the lock of the second OMC. We think it could be related to DARM handoff from B1s2 to B1: maybe, besides the "b1_pd_balance" there is some other value to be changed accordingly, but we are not sure.

SIDE notes: during the shift the ITF unlocked many times because of:

- fast unlocks

- PR misalignment

- earthquake, which excited the suspensions for about one hour

Since we couldn't reach LN3 we couldn't perform the scheduled negative DARM offset scan.

Comments to this report:
mwas - 8:16 Wednesday 23 January 2019 (44487) Print this report
This morning between 6UTC and 7UTC I have made a few more trials of DC read-out.

First changed the mixing of B1s2 and B1_PD2 for the OMC1 error signal (the SDB1_B1x channels), as the ratio of power has now changed from 1/6 to 1/12.
This didn't have an impact.

Then noticed that the when OMC2 goes through the resonance the total power between B1s2 and B1 is not kept constant.
This is because SDB2_B1_DC has been for the past year the average of the two B1 PD powers, and with the current balance weight of 0.0 for B1_PD1 and 1.0 for B1_PD2 it actually corresponds to the total power on the B1 beam (so double the power in the situation with the 50/50 beam splitter).
Tried to divide the darm_b1dc_weight by a factor 2, it kept the total B1s2 and B1 power constant, but the ITF still unlocks.
To make it a bit cleaner (in case SDB2_B1_DC is used somewhere else), changed the value of the SDB2_ONE constant from 1.0 to 0.5, in this way SDB2_B1_DC = 0*B1_PD1 + 0.5*B1_PD2.
This didn't solve the problem either.

Figure 1 summarizes the situation when OMC2 goes through the resonance. The total B1s2 and B1 power is constant 12*B1s2 + B1_PD2 = 18mW.
But the B1p_56MHz_I signal is clearly moving away from zero as if the dark fringe offset is increasing.

An explanation could be that the beam is not well aligned on B1_PD1, and that only ~20% of the light reaches the photodiode.
Images attached to this comment
Search
×

Warning

×